CARA Telecon Minutes Date: June 4, 2001 Time: 1:00 PM Attendance: BU: Cornell: CMU: J. Peterson Northwestern: G. Novak R.I.T: S.A.O.: T. Stark, A. Lane, W. Walsh South Pole: C. Martin, C. Kaminski Chicago: J. Carlstrom, A. Harper, N. Odalen, R. Pernic, R. Landsberg, E. Leitch UCB: W. Holzapfel UCSB: Agenda Items Agenda from May 21, 2001 approved. News from the Pole: C. Martin reported that the past few weeks AST/RO has been plugging away. Weather has not been good. Tau in the region of 1 for 2 or 3 days and then with the stormy weather Tau increased. Tau did plummet to a level of 1 after readings as high as 3. B. Reddall had problems with the glycol loop. Three receivers warmed up but all now back on line. Cooled without interrupting observing with the other receivers. Viper wheels freezing up and M. Newcomb has to go out to telescope and move. This has been a problem for the past two weeks. It now seems to have fixed itself. C. Kaminski reported that the AASTO is working well. The ice plug did build up but this has now fallen off and a new ice plug is starting to build - currently at 2-3 inches high. SODAR is plugging along. Cryogen - the 3000 gallon dewar was weighted. Preliminary numbers show that helium amounts agree with liquid helium level. Still checking with Raytheon on exact expected weights after modifications were made to the dewar. Additional Instrument Reports: DASI - Even after the glycol loop was lost and the 3 receivers warmed, DASI never stopped taking data with the other receivers. News from the Chicago DASI team - J. Kovac gave DASI polarization talk last week. Team will be working on polarization level of ground contamination after next week. So far appears to be a very good season. AST/RO - Preliminary data reduction by R. Chamberlin show 200 micron window. Transmission about 10%; the 350 tipper tau a little above 1.0 when the FTS data was taken. In answer to comment from J. Carlstrom whether the FTS should only be observing when 200 micron window is open (due to low levels of helium at Pole) and how long would a window stay open at Pole. T. Stark stated that FTS takes about 3 hours to set up for observing. The transmission at lower levels can be good for days. ACBAR - W. Holzapfel commented that the drive seems to have fixed itself and now can't get the drive to stick. Power spectrum looks good. Instrument is taking more drift scans and then should go to deep pointing observations on selected blank fields. *** W. Holzapfel, J. Ruhl, G. Novak and J. Peterson need to discuss observing strategy for rest of season. Should SPARO be installed for some observing? W. Holzapfel stated he would set up telecon for this week. J. Peterson has requested that M. Newcomb change wheel bearings, but not if telescope is working. C. Kaminski will check G. Novak email from 2 days ago and make sure Charlie provided all of Giles's questions. CARA - Management Issues: J. Carlstrom reminded the group that he had sent and email to E. Chang and C. Erb requesting CARA Polar Winterover funding. D. Peacock has emailed J. Carlstrom that NSF would supplement the DASI continuation award for the funds required to support a winterover for 2/02-12/02. J. Carlstrom questioned the group if this was acceptable to the other PIs - no exceptions were noted, several positive responses were received by email. ** Carlstrom to reply to Peacock that plan is okay. W. Holzapfel questioned what would be covered by Polar Operations. NSF at 85% of the requested amount funded the ACBAR continuation proposal. W. Holzapfel has been requested by D. Peacock to prepare a revised budget and W. Holzapfel does not want any duplicated effort to be accounted for the in ACBAR budget. I.e. people, supplies, assistance at Pole. ***W. Holzapfel will revise ACBAR budget with a supplemental list on what else is needed to be funded. R. Pernic reported that he has not received any feedback on the CARA equipment list that was submitted to Raytheon. Raytheon still has not set a date for the planning meeting. P. Sullivan is now assisting E. Jensen. J. Carlstrom wants comments from group so he can be armed with the justification on what is needed to support the efforts at Pole. Currently NSF feels that only $200k $300k are required to support CARA Polar Operations when actual numbers have been in the range between $550k-$900k R. Pernic noted that Raytheon had sent out medical kits to all individuals on the SIPs. Since a planning meeting has not been held and we are unsure of the population cap all individual that may have received medical kits may not go to Pole this year. Therefore, PIs should not have their people spend money on medical exams until they know they will be going to Pole. *** R. Pernic to send out email to all PIs explaining the need to hold up on the medical exams.